# AboutMe: Using Self-Descriptions in Webpages to Document the Effects of English Pretraining Data Filters

### v1

@@ 4.1 Topical interests
Similarities in how data filters score topical clusters cut across quality and English filters (Table 5, Table 4). Pairwise correlations of topics’ average English scores across all four langID systems demonstrate high consensus (mean $r_{s}$ = 0.874, SD = 0.038, all $p<0.001$). Surprisingly, Wikipedia perplexity also behaves like fastText langID ($r_{s}$ = 0.860, $p<0.001$). We qualitatively examine 20 random pages from highly filtered clusters, such as fashion, women and online, store, which are often shopping websites. We find that pages with lower English scores list product names or specifications of individual products, and their original content may have been highly visual. [^2] Indeed, manual inspection of current versions of these websites, when available, supports this claim. Indeed, further down the list of commonly highly filtered topical interests are clusters related to photography and art (Appendix C.2). Thus, though text-based LLMs may intend to be comprehensive in knowledge, they exclude information that is primarily communicated via other forms of media.
@@
Surprisingly, we find that across all quality and English langID filters, web content created by individuals is widely preferred (Appendix D.2). For example, Gopher removes 25.9% of webpages by individuals, and 28.3% of those by organizations. One reason for this pattern is that topics that receive overall low scores by data filters are dominated by organizations, such as businesses in the topical clusters products, quality and online, store. However, even when topics are fixed, organizations are still more likely to be removed by quality and langID filters, with only WikiRefs as an exception (Appendix D.2). Webpages by organizations tend to be shorter than those by individuals across and within topics, and their webpages include more non-alphabetic “words”, more repetition, and fewer words from Gopher’s required list (all $p<0.001$, Mann-Whitney $U$-test). [^3] Gopher’s required wordlist includes the, be, to, of, and, that, have, with. See Appendix B for details.
@@ 4.3 Social roles
[…]
The degree to which pages’ self-identified roles are expressed through their text affects filtering scores as well (Figure 4). Within each role’s collection of webpages, we calculate the proportion of a page that contains vocabulary specific to that role. Following past work Zhang et al. (2017); Lucy et al. (2023), we identify role-specific vocabulary using a metric of association between word types and roles, where their normalized pointwise mutual information score is greater than 0.1. We compare how filters score all pages within a role, and how they score a subset of the role’s pages that contain more role-specific words than average. We find that roles that are generally favored by a filter tend to be favored even more if they contain more role-specific words, in contrast to roles that are scored lowest by a filter, which do not benefit from role-specific word use or are penalized further. Exceptions to this pattern are CLD2 and langdetect, two langID filters that score the vast majority (>90%) of pages similarly. These findings suggest that caution may be needed when using pretrained LLMs out-of-the-box for tasks and applications that involve language specific to some domains.
@@ 4.4 Geography
However, geographic filtering patterns are not only explained by topical differences. As expected, most English filters prefer subregions with “core anglophone” countries: Northern America (Canada and US), Northern Europe (UK), and Australia & New Zealand (Figure 5). Subregions with lower English document-level scores contain more non-English paragraphs across all langID systems (mean $r$ = -0.791, SD = 0.084, all $p<0.05$). By examining these “non-English” paragraphs, we observe two reasons for why a page may not be “English” enough. First, langID can mislabel English text Caswell et al. (2020), such as content containing names of products, people, and non-anglophone locations. Second, some web pages are indeed multilingual, either code-switching or including multiple translations of the same content. LangID is usually applied at the document-level during data curation, and some systems may assume monolingual inputs Zhang et al. (2018). Non-English paragraphs in AboutMe reflect their geography, e.g. Chinese in Eastern Asia, Spanish in Latin America, and Polish in Eastern Europe. Thus, simply choosing to communicate in English is not necessarily grounds for inclusion, and how English is situated across and within webpages matters.
Finally, we examine how filtering of geographic locations may relate to their relative global status. Past work has suggested that some NLP models may favor wealthier countries Zhou et al. (2022). In our case, we do not find a significant relationship between a country’s filter scores and their gross domestic product (Appendix G.3).
@@ 4.5 Regression analysis
Finally, we investigate the relative importance of various dimensions highlighted in previous sections for model-based filters. That is, what matters more during data filtering: who you are, or where you are from?
@@
[…]
@@ 8 Acknowledgements
We thank Nicholas Tomlin, Naitian Zhou, and Kyle Lo for helpful conversations and feedback.
@@ B.1 Filter reproduction
[…]
@@ C.2 Additional filtering results
[…]
@@ D.2 Additional filtering results
[…]
@@
[…]
@@ Appendix F Social roles
[…]
@@ F.2 Token classification
Our labeled spans are whole words, but RoBERTa sometimes labels parts of words. When we run inference on all individuals’ about pages, we find that it nearly always tags all wordpieces in a positive span correctly. Still, 3.3% of tagged words are partially tagged, e.g. play-mate, trend-set-ter, mom-my. From manual inspection of these cases, it seems like partially tagged words are usually social roles. Thus, we evaluate at the word-level, and count words as social roles if any of its wordpieces is tagged as one. Table 16 shows common terms extracted from all individuals’ about pages. Some tagged words are part of hyphenated phrases, e.g. co-president. We do not consider common prefixes (e.g. vice-, ex-) and suffixes (e.g. -elect, -in-law) as individual roles during analysis.
@@ F.3 Occupation hierarchy
Arts, Design, Entertainment, Sports, & Media
1.1M
artist, director, designer, writer, photographer, musician, player
Production
620K
designer, engineer, maker, builder, operator, mechanic
Community & Social Service
452K
therapist, educator, advisor, pastor, activist, social worker
Computer & Mathematical
365K
engineer, developer, scientist, strategist, programmer
Educational Instruction & Library
308K
teacher, professor, lecturer, curator, tutor, graduate student
Healthcare Practitioners and Technical
300K
therapist, nurse, doctor, nutritionist, surgeon, midwife
Management
291K
president, manager, dean, administrator, medical director
Architecture and Engineering
288K
architect, technician, electrical engineer, technologist, tester
Business and Financial Operations
250K
analyst, accountant, marketer, investor, management consultant
Personal Care and Service
205K
trainer, yoga teacher, stylist, makeup artist, caregiver
@@
Figure 10: For some filters, we find statistically significant relationships between an occupation’s prestige or salary ($y$-axes) and average filtering scores ($x$-axes).
@@ F.4 Additional filtering results
CLD3 -0.1302 -0.0339
langdetect -0.1353 -0.0833
[…]
Wiki -0.1051 -0.2336***
WikiRefs -0.1076 -0.1713**
OpenWeb -0.1349 -0.2280***
WikiWebBooks -0.1115 -0.2238***
Gopher -0.2071*** -0.1139
[…]
@@
[…]
@@ G.3 Additional filtering results
[…]
@@ Appendix H Regression
[…]
@@
Intercept ${}^{***}$
Topic: news, media, content  ${}^{***}$
Topic: film, production, festival  ${}^{***}$
Topic: writing, books, book  ${}^{***}$
Topic: music, band, musical  ${}^{***}$
Topic: research, university, science  ${}^{***}$
Topic: service, cleaning, repair ${}^{***}$
Topic: hair, beauty, skin ${}^{***}$
Topic: insurance, care, dental ${}^{***}$
Topic: home, homes, family ${}^{***}$
Topic: estate, real, property ${}^{***}$
Region: Americas  ${}^{***}$
Region: Asia  0.002
Region: Europe  ${}^{***}$
Region: Oceania  ${}^{***}$
Individual  ${}^{***}$
Core anglophone ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: news, media, content  ${}^{***}$
Topic: people, world, work  ${}^{***}$
Topic: software, data, development  ${}^{***}$
Topic: writing, books, book  ${}^{***}$
Topic: like, love, time  ${}^{***}$
Topic: service, cleaning, repair ${}^{***}$
Topic: estate, real, property ${}^{***}$
Topic: quality, equipment, production ${}^{***}$
Topic: home, homes, family ${}^{***}$
Topic: insurance, care, dental ${}^{***}$
Region: Americas  ${}^{***}$
Region: Asia -0.003
Region: Europe  ${}^{***}$
Region: Oceania  ${}^{***}$
Individual  ${}^{***}$
Core anglophone ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: news, media, content  ${}^{***}$
Topic: club, members, association  ${}^{***}$
Topic: music, band, musical  ${}^{***}$
Topic: film, production, festival  ${}^{***}$
Topic: research, university, science  ${}^{***}$
Topic: service, cleaning, repair ${}^{***}$
Topic: hair, beauty, skin ${}^{***}$
Topic: home, homes, family ${}^{***}$
Topic: furniture, jewelry, quality ${}^{***}$
Topic: products, quality, product ${}^{***}$
Region: Americas $*{}^{*\phantom{**}}$
Region: Asia ${}^{***}$
Region: Europe  ${}^{***}$
Region: Oceania ${}^{***}$
Individual ${}^{***}$
Core anglophone ${}^{***}$
[…]
@@
Intercept  ${}^{***}$
Topic: film, production, festival  ${}^{***}$
Topic: music, band, musical  ${}^{***}$
Topic: research, university, science  ${}^{***}$
Topic: club, members, association -0.007
Topic: news, media, content ${}^{***}$
Topic: blog, like, love ${}^{***}$
Topic: service, cleaning, repair ${}^{***}$
Topic: hair, beauty, skin ${}^{***}$
Topic: life, yoga, help ${}^{***}$
Topic: like, love, time ${}^{***}$
Region: Americas  ${}^{***}$
Region: Asia  ${}^{***}$
Region: Europe  ${}^{***}$
Region: Oceania  ${}^{***}$
Individual  ${}^{***}$
Core anglophone ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: law, legal, firm  ${}^{***}$
Topic: god, church, christ  ${}^{***}$
Topic: insurance, care, dental  ${}^{***}$
Topic: research, university, science  ${}^{***}$
Topic: financial, clients, investment  ${}^{***}$
Topic: online, store, shopping ${}^{***}$
Topic: fashion, women, brand ${}^{***}$
Topic: products, quality, product ${}^{***}$
Topic: quality, equipment, production ${}^{***}$
Topic: furniture, jewelry, quality ${}^{***}$
Region: Americas  ${}^{***}$
Region: Asia ${}^{***}$
Region: Europe  ${}^{***}$
Region: Oceania  ${}^{***}$
Individual  ${}^{***}$
Core anglophone  0.002
[…]
@@
[…]
@@
Intercept ${}^{***}$
Topic: law, legal, firm  ${}^{***}$
Topic: insurance, care, dental  ${}^{***}$
Topic: children, child, school  ${}^{***}$
Topic: god, church, christ  ${}^{***}$
Topic: financial, clients, investment  ${}^{***}$
Topic: online, store, shopping ${}^{***}$
Topic: quality, equipment, production ${}^{***}$
Topic: fashion, women, brand ${}^{***}$
Topic: products, quality, product ${}^{***}$
Topic: com, www, https ${}^{***}$
Region: Americas ${}^{***}$
Region: Asia ${}^{***}$
Region: Europe -0.004
Region: Oceania ${}^{***}$
Individual  ${}^{***}$
Core anglophone  ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: solutions, technology, business  ${}^{***}$
Topic: marketing, digital, media  ${}^{***}$
Topic: insurance, care, dental  ${}^{***}$
Topic: financial, clients, investment  ${}^{***}$
Topic: services, service, clients  ${}^{***}$
Topic: quality, equipment, production ${}^{***}$
Topic: com, www, https ${}^{***}$
Topic: fashion, women, brand -0.001
Topic: company, products, quality -0.001
Topic: online, store, shopping  0.005
Region: Americas ${}^{***}$
Region: Asia ${}^{***}$
Region: Europe ${}^{***}$
Region: Oceania ${}^{***}$
Individual  ${}^{***}$
Core anglophone  ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: solutions, technology, business  ${}^{***}$
Topic: insurance, care, dental  ${}^{***}$
Topic: services, service, clients  ${}^{***}$
Topic: service, cleaning, repair  ${}^{***}$
Topic: financial, clients, investment  ${}^{***}$
Topic: online, store, shopping ${}^{***}$
Topic: quality, equipment, production ${}^{***}$
Topic: fashion, women, brand ${}^{***}$
Topic: com, www, https ${}^{***}$
Topic: music, band, musical ${}^{***}$
Region: Americas ${}^{***}$
Region: Asia ${}^{***}$
Region: Europe -0.002
Region: Oceania ${}^{***}$
Individual  ${}^{***}$
Core anglophone  ${}^{***}$
[…]
@@
Intercept ${}^{***}$
Topic: solutions, technology, business  ${}^{***}$
Topic: construction, project, projects  ${}^{***}$
Topic: services, service, clients  ${}^{***}$
Topic: children, child, school  ${}^{***}$
Topic: marketing, digital, media  ${}^{***}$
Topic: online, store, shopping ${}^{***}$
Topic: fashion, women, brand ${}^{***}$
Topic: car, vehicle, auto ${}^{***}$
Topic: com, www, https ${}^{***}$
Topic: products, quality, product ${}^{***}$
Region: Americas ${}^{***}$
Region: Asia ${}^{***}$
Region: Europe ${}^{***}$
Region: Oceania ${}^{***}$
Individual  ${}^{***}$
Core anglophone  ${}^{***}$
[…]

### v3

@@ 4.1 Topical interests
Similarities in how data filters score topical clusters cut across quality and English filters (Table 4, Table 5). Pairwise correlations of topics’ average English scores across all four langID systems have high consensus (mean $r_{s}$ = 0.874, SD = 0.038, all $p<0.001$). Surprisingly, Wikipedia perplexity also behaves like fastText langID ($r_{s}$ = 0.860, $p<0.001$). We qualitatively examine 20 random pages from highly filtered clusters, e.g. fashion, women and online, store. We find that pages with lower English scores list product names or specifications of individual products, and their original content may have been highly visual. [^2] Indeed, manual inspection of current versions of these websites, when available, supports this claim. Indeed, further down the list of commonly highly filtered topical interests are clusters related to photography and art (Appendix C.2). Thus, though text-based LLMs may intend to be comprehensive in knowledge, they exclude information that is primarily communicated via other forms of media.
@@
Surprisingly, we find that across all quality and English langID filters, web content created by individuals is widely preferred (Appendix D.2). For example, Gopher removes 25.9% of webpages by individuals, and 28.3% of those by organizations. One reason for this pattern is that topics that receive overall low scores by data filters are dominated by organizations, e.g. businesses in the clusters products, quality and online, store. However, even when topics are fixed, organizations are still more likely to be removed by nearly all filters, with only WikiRefs as an exception (Appendix D.2). Webpages by organizations tend to be shorter than those by individuals across and within topics, and their webpages include more non-alphabetic “words”, more repetition, and fewer words from Gopher’s required list (all $p<0.001$, Mann-Whitney $U$-test). [^3] Gopher’s required wordlist includes the, be, to, of, and, that, have, with. See Appendix B for details.
@@ 4.3 Social roles
Filters’ social role preferences mirror those for topical interests (Table 6), such as programming occupations and software, data being preferred by OpenWeb, and correspondent matching the often “high quality” topic of news, media. We measure the relationship between model-based filter scores and two metrics that reflect occupations’ societal status: their O*NET salary estimate and Hughes et al. (2022)’s survey-based ratings of prestige. We find small yet significant relationships between occupational prestige and model-based quality scores ($p<0.001$, Appendix F.4). That is, pages linked to lower prestige occupations are filtered more by WikiWebBooks, OpenWeb, WikiRefs, Wiki, and Wikippl.
The degree to which pages’ self-identified roles are expressed through their text affects filtering as well (Figure 4). Within each role’s collection of webpages, we calculate the proportion of each page that contains vocabulary specific to that role. Following past work Zhang et al. (2017); Lucy et al. (2023), we identify role-specific vocabulary using a metric of association between word types and roles, where their normalized pointwise mutual information (NPMI) score is greater than 0.1. We compare how filters score all pages within a role, and how they score a subset of the role’s pages that contain more role-specific words than average. We find that roles that are generally favored by a filter tend to be favored even more when their pages contain more role-specific words, in contrast to roles that are scored lowest by a filter, which do not benefit from role-specific word use or are penalized further. Exceptions to this pattern are CLD2 and langdetect, two langID filters that score the vast majority (>90%) of pages similarly. These findings suggest that caution may be needed when using pretrained LLMs out-of-the-box for tasks and applications that involve language specific to some domains.
@@ 4.4 Geography
However, geographic filtering patterns are not only explained by topical differences. As expected, most English filters prefer subregions with “core anglophone” countries: Northern America (Canada and US), Northern Europe (UK), and Australia & New Zealand (Figure 5). Subregions with lower English document-level scores contain more non-English paragraphs across all langID systems (mean $r$ = -0.791, SD = 0.084, all $p<0.05$). By examining these “non-English” paragraphs, we observe two reasons for why a page may not be “English” enough. First, langID can mislabel English text Caswell et al. (2020), such as content containing names of products, people, and non-anglophone locations. Second, some web pages are indeed multilingual, either code-switching or including multiple translations of the same content. LangID is usually applied at the document-level during data curation, and some systems may assume monolingual inputs Zhang et al. (2018). Non-English paragraphs in AboutMe reflect their geography, e.g. Chinese in Eastern Asia, Spanish in Latin America, and Polish in Eastern Europe. Thus, simply choosing to communicate in English is not necessarily grounds for inclusion, and how English is situated within webpages matters.
In addition, we examine how filtering of geographic locations may relate to their relative global status. Past work has suggested that some NLP models may favor wealthier countries Zhou et al. (2022). In our case, we do not find a significant relationship between a country’s filter scores and their gross domestic product (Appendix G.3).
[…]
@@ 4.5 Regression analysis
Finally, we investigate the relative importance of the social dimensions highlighted in previous sections for model-based filters. That is, what matters more: who you are, or where you are from?
We run several ordinary least squares regressions with different filters’ scores as dependent variables, and topical interests, continental regions, individual/organization status, core anglophone status, and pages’ character length as independent variables. All independent variables, except for length, are binary or categorical. We $z$-score standardize each regression’s dependent variable so that coefficients are similarly interpretable across them. We find that length has a positive effect ($p<0.001$) on all model-based filter scores except for Wikippl, and many topical variables have stronger effects on filter scores than other variable types, especially among quality classifiers (Figure 6, Appendix H). Though earlier we noted that some subregions’ filtering scores may be due to topical skew (§4.4), even when controlling for topic, Asia still has the most negative coefficients relative to other continental regions for all langID filters.
4.6 Summary
We find shared patterns in how quality and English filters score websites delineated by social aspects such as social roles and geography. Differences in how quality filters behave depend on both how they are implemented and their choice of “high quality” reference corpora (§4.1). This latter factor results in notions of “quality” being associated with certain topical domains e.g. news and media, and these topical preferences then lead to filters privileging content specific to some social roles over others’ (§4.3). Finally, common langID classifiers can overlook English content in non-anglophone regions of the world, especially Asia, even when controlling for other variables such as webpage length and topic (§4.5).
5 Related Work
@@
Though a nascent research area, analyses of pretraining data curation decisions have also been the focus of other recent literature Longpre et al. (2023); Soldaini et al. (2024). For example, Gao (2021) showed that discarding too much pretraining data using the Pile’s quality filter can lead to worse downstream task performance. Our work is closest in spirit to Gururangan et al. (2022), who use a dataset of high school newspapers to show that text from wealthier, more educated, and urban areas are more likely to be considered high quality by GPT-3’s model-based quality filter. Similarly, concurrent work by Hong et al. (2024) found that image-text CLIP-filtering for visual language models excludes data from LGBTQ+ people, older women, and younger men at higher rates. We also critically examine social aspects of quality filtering at scale, but across a range of text filters.
@@ 9 Acknowledgements
We thank Nicholas Tomlin, Naitian Zhou, and Kyle Lo for helpful conversations and feedback. We also thank our anonymous reviewers for their thoughtful reviews. This work was supported in part by the National Science Foundation (IIS-1942591).
@@ B.1 Filter reproduction
All model-based quality filters, except for Wikippl, use the same implementation and parameter choices as the reproduction of GPT-3’s quality filter by Gururangan et al. (2022).
@@ C.2 Additional filtering results
Quality: Wiki Quality: Wikippl English: fastText
@@ D.2 Additional filtering results
Quality: Wiki Quality: Wikippl English: fastText
@@
Table 13: The percentage of topics where individuals have significantly higher model-based filter scores on average than organizations in the same topic (Mann Whitney $U$-test $p<0.001$). Note that for Wikippl, we reverse perplexity scores so that the higher, the better, to match the same direction as other model-based filters.
@@ Appendix F Social roles
Quality: Wiki Quality: Wikippl English: fastText
@@ F.2 Token classification
Our labeled spans are whole words, but RoBERTa sometimes labels parts of words. When we run inference on all individuals’ about pages, we find that it nearly always tags all wordpieces in a positive span correctly. Still, 3.3% of tagged words are partially tagged, e.g. play-mate, trend-set-ter, mom-my. From manual inspection of these cases, it seems like partially tagged words are usually social roles. Thus, we evaluate at the word-level, and count words as social roles if any of its wordpieces is tagged as one.
Table 16 shows common terms extracted from all individuals’ about pages. Some tagged words are part of hyphenated phrases, e.g. co-president. We do not consider common prefixes (e.g. vice-, ex-) and suffixes (e.g. -elect, -in-law) as individual roles during analysis.
@@ F.3 Occupation hierarchy
Arts, Design, Entertainment, Sports, & Media 1.1M artist, director, designer, writer, photographer, musician, player
Production 620K designer, engineer, maker, builder, operator, mechanic
Community & Social Service 452K therapist, educator, advisor, pastor, activist, social worker
Computer & Mathematical 365K engineer, developer, scientist, strategist, programmer
Educational Instruction & Library 308K teacher, professor, lecturer, curator, tutor, graduate student
Healthcare Practitioners and Technical 300K therapist, nurse, doctor, nutritionist, surgeon, midwife
Management 291K president, manager, dean, administrator, medical director
Architecture and Engineering 288K architect, technician, electrical engineer, technologist, tester
Business and Financial Operations 250K analyst, accountant, marketer, investor, management consultant
Personal Care and Service 205K trainer, yoga teacher, stylist, makeup artist, caregiver
@@
Figure 10: For some filters, we find statistically significant relationships between an occupation’s prestige or salary ($y$-axes) and average filtering scores ($x$-axes), $p<0.001$.
@@ F.4 Additional filtering results
CLD3 0.1302 0.0339
langdetect 0.1353 0.0833
Wikippl 0.2102*** 0.2176***
Wiki 0.1051 0.2336***
WikiRefs 0.1076 0.1713**
OpenWeb 0.1349 0.2280***
WikiWebBooks 0.1115 0.2238***
Gopher 0.2071*** 0.1139
Table 18: Pearson correlation values between prestige or salary and filters’ scores (higher = less filtered). For Gopher, we use the negated rate of page removal as the “score” since that filter does not output a single numerical score. We also negate scores for Wikippl so that its values can be interpreted similarly to other rows, since higher perplexities values get filtered more rather than less. Significance: *$p<0.05$, **$p<0.01$, ***$p<0.001$, with Bonferroni correction for 20 comparisons.
@@
We find that for salary, two filters (Wikippl and Gopher) show statistically significant relationships with salary, where higher-paid occupations are filtered less (Table 18, Figure 10). For prestige, all quality filters except for Gopher show a statistically significant relationship. The most and least filtered social roles shown in Table 14 intuitively reflect these trends. For example, tech-related engineering occupations are highly scored by OpenWeb, and these tend to have prestige scores over 60.
@@ G.3 Additional filtering results
We limit country-level filtering analyses only to countries that appear at least 500 times in our dataset, to ensure the patterns we find are over enough samples. We find weak Pearson correlations ($p<0.05$, with Bonferroni correction) between a country’s GDP and their average filtering scores for fastText, CLD3, and Wikippl. However, these results are only due to a single outlier, China, which is often the most filtered country but also very high in GDP. After removing this outlier, all $p$ values are insignificant, and thus we do not confidently conclude any broad relationship between wealth status and filtering.
@@ Appendix H Regression
In §4.5, we run nine ordinary least squares regressions, one for each model-based filter, to investigate how different aspects of websites that we extract relate to filtering scores. To transform categorical variables into dummy binary variables, we use Africa as the base category for region, and art, gallery as the base category for topical interests. Since the directionality of how Wikippl should be interpreted is the opposite of other filters’ scores, we negate its scores before performing its regression. Tables 21-29 show the results of these regressions in more detail. For clarity of interpretation, we include coefficients for only a subset of all topics with the most positive and negative effects in each regression. The topics with highest and lowest coefficients tend to reflect ones that are highly retained or removed by a filter.
@@
Intercept -1.170∗∗∗
Topic: news, media, content  0.414∗∗∗
Topic: film, production, festival  0.319∗∗∗
Topic: writing, books, book  0.206∗∗∗
Topic: music, band, musical  0.176∗∗∗
Topic: research, university, science  0.152∗∗∗
Topic: service, cleaning, repair -0.567∗∗∗
Topic: hair, beauty, skin -0.524∗∗∗
Topic: insurance, care, dental -0.504∗∗∗
Topic: home, homes, family -0.499∗∗∗
Topic: estate, real, property -0.476∗∗∗
Region: Americas  0.090∗∗∗
Region: Asia  ${}^{\phantom{***}}$
Region: Europe  0.078∗∗∗
Region: Oceania  0.074∗∗∗
Individual  0.123∗∗∗
Core anglophone -0.147∗∗∗
log2(# of characters)  0.142∗∗∗
@@
Intercept -0.803∗∗∗
Topic: news, media, content  0.772∗∗∗
Topic: people, world, work  0.359∗∗∗
Topic: software, data, development  0.315∗∗∗
Topic: writing, books, book  0.315∗∗∗
Topic: like, love, time  0.255∗∗∗
Topic: service, cleaning, repair -0.387∗∗∗
Topic: estate, real, property -0.385∗∗∗
Topic: quality, equipment, production -0.353∗∗∗
Topic: home, homes, family -0.342∗∗∗
Topic: insurance, care, dental -0.302∗∗∗
Region: Americas  0.104∗∗∗
Region: Asia ${}^{\phantom{***}}$
Region: Europe  0.089∗∗∗
Region: Oceania  0.063∗∗∗
Individual  0.088∗∗∗
Core anglophone -0.072∗∗∗
log2(# of characters)  0.080∗∗∗
@@
Intercept -0.917∗∗∗
Topic: news, media, content  0.490∗∗∗
Topic: club, members, association  0.363∗∗∗
Topic: music, band, musical  0.347∗∗∗
Topic: film, production, festival  0.329∗∗∗
Topic: research, university, science  0.253∗∗∗
Topic: service, cleaning, repair -0.546∗∗∗
Topic: hair, beauty, skin -0.46∗∗∗
Topic: home, homes, family -0.414∗∗∗
Topic: furniture, jewelry, quality -0.411∗∗∗
Topic: products, quality, product -0.406∗∗∗
Region: Americas ${}^{*\phantom{**}}$
Region: Asia -0.029∗∗∗
Region: Europe  0.019∗∗∗
Region: Oceania -0.010∗∗∗
Individual -0.031∗∗∗
Core anglophone -0.049∗∗∗
log2(# of characters)  0.114∗∗∗
@@
Intercept  0.186∗∗∗
Topic: film, production, festival  0.181∗∗∗
Topic: music, band, musical  0.117∗∗∗
Topic: research, university, science  0.112∗∗∗
Topic: club, members, association ${}^{\phantom{***}}$
Topic: news, media, content -0.097∗∗∗
Topic: blog, like, love -0.487∗∗∗
Topic: service, cleaning, repair -0.468∗∗∗
Topic: hair, beauty, skin -0.466∗∗∗
Topic: life, yoga, help -0.447∗∗∗
Topic: like, love, time -0.443∗∗∗
Region: Americas  0.029∗∗∗
Region: Asia  0.019∗∗∗
Region: Europe  0.041∗∗∗
Region: Oceania  0.039∗∗∗
Individual  0.056∗∗∗
Core anglophone -0.165∗∗∗
log2(# of characters)  0.016∗∗∗
@@
Intercept -1.316∗∗∗
Topic: law, legal, firm  0.121∗∗∗
Topic: god, church, christ  0.118∗∗∗
Topic: insurance, care, dental  0.075∗∗∗
Topic: research, university, science  0.071∗∗∗
Topic: financial, clients, investment  0.07∗∗∗
Topic: online, store, shopping -0.38∗∗∗
Topic: fashion, women, brand -0.375∗∗∗
Topic: products, quality, product -0.341∗∗∗
Topic: quality, equipment, production -0.296∗∗∗
Topic: furniture, jewelry, quality -0.288∗∗∗
Region: Americas  0.047∗∗∗
Region: Asia -0.102∗∗∗
Region: Europe  0.063∗∗∗
Region: Oceania  0.015∗∗∗
Individual  0.041∗∗∗
Core anglophone  ${}^{\phantom{***}}$
log2(# of characters)  0.132∗∗∗
@@
Table 25: Regression results for the quality filter Wikippl. *$p<0.05$, **$p<0.01$, and ***$p<0.001$.
@@
Intercept -2.154∗∗∗
Topic: law, legal, firm  0.310∗∗∗
Topic: insurance, care, dental  0.292∗∗∗
Topic: children, child, school  0.282∗∗∗
Topic: god, church, christ  0.276∗∗∗
Topic: financial, clients, investment  0.264∗∗∗
Topic: online, store, shopping -0.37∗∗∗
Topic: quality, equipment, production -0.293∗∗∗
Topic: fashion, women, brand -0.262∗∗∗
Topic: products, quality, product -0.239∗∗∗
Topic: com, www, https -0.177∗∗∗
Region: Americas -0.078∗∗∗
Region: Asia -0.152∗∗∗
Region: Europe ${}^{\phantom{***}}$
Region: Oceania -0.057∗∗∗
Individual  0.073∗∗∗
Core anglophone  0.112∗∗∗
log2(# of characters)  0.206∗∗∗
@@
Intercept -0.211∗∗∗
Topic: solutions, technology, business  0.121∗∗∗
Topic: marketing, digital, media  0.098∗∗∗
Topic: insurance, care, dental  0.098∗∗∗
Topic: financial, clients, investment  0.096∗∗∗
Topic: services, service, clients  0.094∗∗∗
Topic: quality, equipment, production -0.076∗∗∗
Topic: com, www, https -0.035∗∗∗
Topic: fashion, women, brand ${}^{\phantom{***}}$
Topic: company, products, quality ${}^{\phantom{***}}$
Topic: online, store, shopping  ${}^{\phantom{***}}$
Region: Americas -0.047∗∗∗
Region: Asia -0.164∗∗∗
Region: Europe -0.050∗∗∗
Region: Oceania -0.059∗∗∗
Individual  0.011∗∗∗
Core anglophone  0.098∗∗∗
log2(# of characters)  0.015∗∗∗
@@
Intercept -1.330∗∗∗
Topic: solutions, technology, business  0.175∗∗∗
Topic: insurance, care, dental  0.158∗∗∗
Topic: services, service, clients  0.153∗∗∗
Topic: service, cleaning, repair  0.148∗∗∗
Topic: financial, clients, investment  0.146∗∗∗
Topic: online, store, shopping -0.081∗∗∗
Topic: quality, equipment, production -0.075∗∗∗
Topic: fashion, women, brand -0.053∗∗∗
Topic: com, www, https -0.041∗∗∗
Topic: music, band, musical -0.012∗∗∗
Region: Americas -0.021∗∗∗
Region: Asia -0.174∗∗∗
Region: Europe ${}^{\phantom{***}}$
Region: Oceania -0.033∗∗∗
Individual  0.013∗∗∗
Core anglophone  0.069∗∗∗
log2(# of characters)  0.125∗∗∗
@@
Intercept -0.560∗∗∗
Topic: solutions, technology, business  0.042∗∗∗
Topic: construction, project, projects  0.034∗∗∗
Topic: services, service, clients  0.031∗∗∗
Topic: children, child, school  0.029∗∗∗
Topic: marketing, digital, media  0.028∗∗∗
Topic: online, store, shopping -0.069∗∗∗
Topic: fashion, women, brand -0.061∗∗∗
Topic: car, vehicle, auto -0.038∗∗∗
Topic: com, www, https -0.037∗∗∗
Topic: products, quality, product -0.031∗∗∗
Region: Americas -0.025∗∗∗
Region: Asia -0.071∗∗∗
Region: Europe -0.027∗∗∗
Region: Oceania -0.038∗∗∗
Individual  0.014∗∗∗
Core anglophone  0.053∗∗∗
log2(# of characters)  0.055∗∗∗